OpenStudy (badsheep):

A motorcyclist changes the velocity of his bike from 20.0 meters/second to 35.0 meters/second under a constant acceleration of 4.00 meters/second2. How long does it take the bike to reach the final velocity?

OpenStudy (whpalmer4):

this is a physics question, and should have been posted there instead of in chemistry. for constant acceleration, \[v_f - v_i = at\] you know \(v_f = 35.0 \text{ m/s}\), \(v_i = 20.0 \text{ m/s}\), \(a = 4.00 \text{ m/s}^2\) Solve equation for \(t\) and plug in the numbers...

OpenStudy (badsheep):

its chemistry

OpenStudy (badsheep):

well on chemistry any way